Assume the sales are predicted to be 3309contribution margin is 1115 in a net loss of 200 is anticipated the break even point in sales
Answer #1

Contribution margin ratio = contribution margin / sales = 1115 / 3309 = 33.70%

Fixed expense = Contribution margin + Net loss

= 1115 + 200 = 1315

Break even point in sales = Fixed expense / Contribution margin ratio

= 1315 / 0.3370 = 3903